Nailaaaaaa:

A chemist mixes 75.0 g of an unknown substance at 96.5°C with 1,150 g of water at 25.0°C. If the final temperature of the system is 37.1°C, what is the specific heat capacity of the substance? Use 4.184 J / g °C for the specific heat capacity of water. I need an explanation to this question not solution...

xxemilianaxx:

Asked and answered here: https://questioncove.com/updates/4fa6f6a0e4b029e9dc366363 Also keep in mind heat lost = heat gained

Nailaaaaaa:

That's just teh solution, I need explanation for this question...

xxemilianaxx:

Ok wait. You need a explaination but to what exactly, The solution to the problem?

Vocaloid:

To elaborate on my solution from the other post q = mC(Tf-Ti) where q is heat, C is specific heat, Tf is final temperature, Ti is final temperature all the heat that was lost from the unknown substance is gained by the water heat lost by the substance q = -75.0(C)(37.1-96.5) heat gained by the water: q = 1150(4.184)(37.1-25) set them equal to each other and solve for C

Nailaaaaaa:

Yes I need a scientific explanation rather than mathematical reasoning or solution

Nailaaaaaa:

"all the heat that was lost from the unknown substance is gained by the water" this is a very good explanation but could you just elaborate and add more explanation to this

Vocaloid:

Energy is conserved. When the substance is added to water, it loses some amount of heat. That same amount of heat is added to the water. Since we know: the mass of the water, the specific heat of the water, and the temperature change of the water, we know how much heat the water gained via q = mc delta T We can then use that heat to solve for the specific heat of the substance

Nailaaaaaa:

thank you